OpenStudy (judson_rockets):

What is the side length of a cube with a volume of 4913 m3? m

OpenStudy (mathmale):

Work backwards. What's the formula for the volume of a cube of side length s? Write out this formula and then set it = to 4913 m^3. Then find the cube root of both sides.

RhondaSommer (rhondasommer):

do you have options? if so; what are they?

OpenStudy (judson_rockets):

I dont

RhondaSommer (rhondasommer):

spo volume of a cube is what

RhondaSommer (rhondasommer):

like; how do you get the volume of a cube

OpenStudy (judson_rockets):

U mutiply

OpenStudy (mathmale):

Judson, would you please Google "volume of cube"? It makes no sense at all to try to solve a problem like this one when you don't know the formula for "volume of a cube." You will need to know this and other basic volume formulas later on.

OpenStudy (judson_rockets):

kk

OpenStudy (judson_rockets):

I dont know how to get the answer im sorry

OpenStudy (dani_rose):

Formula for volume of a cube \[V=a^3\]

OpenStudy (judson_rockets):

So do i divide

RhondaSommer (rhondasommer):

if \[a^3=4913 m^3\]

OpenStudy (mathmale):

Earlier I suggested that y ou find the CUBE ROOT of both sides of this equation. That should give you the side length, a.

OpenStudy (judson_rockets):

I got an 86% on the test

RhondaSommer (rhondasommer):

\[\mathrm{Subtract\:}\left(4913\right)\mathrm{\:from\:both\:sides}\]

RhondaSommer (rhondasommer):

\[a^3-4913=0\] factor \[\:a^3-4913\]\[\left(a-17\right)\left(a^2+17a+17^2\right)\]

RhondaSommer (rhondasommer):

then just do \[\left(a-17\right)=0\]

RhondaSommer (rhondasommer):

\[\left(a-17\right)=0\] add 17 on both sides. a=17

RhondaSommer (rhondasommer):

make sense?